OpenStudy (anonymous):

how do i factor (p-q)(p+q)+(p-q) into (p−q)(p+q+1)?

OpenStudy (jhannybean):

treat \((p-q) =a\) and \((p+q)=b\), it's easier to look at these if they're letters instead of big, scary terms

OpenStudy (jhannybean):

\[(p-q)(p+q)+(p-q) \implies ab+a\] That means if we factor out the common term \(a\) we would have \[=a(b+1)\]

OpenStudy (jhannybean):

Do you understand, @one2345 ?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

\[\large \color{red}{\checkmark}\]

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

Np